A soft drink machine outputs a mean of 2828 ounces per cup. The machine's output is normally distributed with a standard deviation of 22 ounces. What is the probability of filling a cup between 3030 and 3131 ounces? Round your answer to four decimal places.
Answer #1

Solution :

Given that ,

mean = = 28

standard deviation = = 22

P(30 < x < 31) = P[(30 - 28)/ 22) < (x - ) /  < (31 - 28) / 22) ]

= P(0.09 < z < 0.14)

= P(z < 0.14) - P(z < 0.09)

= 0.5557 - 0.5359

= 0.0198

Probability = 0.0198